A block of wood of mass m = 250 g is being pulled along a surface by a force F = 20 N. The force of friction between the block

and the surface has a magnitude of 15 N as

shown. Calculate:

a) the acceleration of the block.

b) the distance moved, starting from rest, in 5.0 s.


1
Expert's answer
2021-12-06T09:42:26-0500

a) The acceleration:


"a=\frac{F_\text{net}}{m}=\frac{F-f}{m}=\frac{20-15}{0.25}=20\text{ m/s}^2."